OpenStudy (anonymous):

Find the discontinuities of f(x) = 2x^2 + 5x - 12/ x+4

OpenStudy (anonymous):

you may not divide by 0, so set the denominator equal to 0 and solve for (x\) more or less in your head

OpenStudy (anonymous):

i meant solve for \(x\)

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

No :o I dont see how I can simplify the numerator. ive done 3 problems But I dont know how to do this one. If someone could just show me how to simplify it and get it in simplest form i would know How to find the discontinuitie/discontinuities of the function

OpenStudy (anonymous):

think of where you could get a 0 in the denominator, check -4 for example

OpenStudy (anonymous):

Wait since it cant be simplified would I plug in -4 for x?

OpenStudy (anonymous):

well negative 4 for x in the numerator?

OpenStudy (anonymous):

the common denominator will be x+4, the place where that = 0 is at x = -4

OpenStudy (anonymous):

thats what i was saying.

OpenStudy (anonymous):

So i plug in -4 for x in the numerator?

OpenStudy (anonymous):

here is the deal you cannot divide by zero aka you cannot have a zero in the denominator so it will be discontinuous if \(x+4=0\) i.e. if \(x=-4\)

OpenStudy (anonymous):

that is all you have to do, there is nothing else

OpenStudy (anonymous):

I'm not sure I understand your question, you don't do anything with the numerator, if you want to convert the equation then multiply each numerator and denominator by (x+4)

OpenStudy (anonymous):

that is, no need to factor the numerator, unless you just feel like doing it

OpenStudy (anonymous):

I mean Like -4 is x for the discontinuous so to find why I would plug in -4 for x to find the y of the discontinous right?

OpenStudy (anonymous):

if you want you can write \[\frac{(2x-3)(x+4)}{x+4}=2x-3\] but it is still discontinuous at \(x=-4\)

OpenStudy (anonymous):

if you plug in \(-4\) for \(x\) you will get \(\frac{0}{0}\)

OpenStudy (anonymous):

the question "find the discontinuities" refers to the \(x\) value unless you have a weird teacher

OpenStudy (anonymous):

now it is true that you can replace \(x\) by \(-4\) in the expression \(2x-3\) and if you do you get \(2\times -4-3=-11\) but the answer to "find the discontinuity" is " it is discontinuous at \(x=-4\)

OpenStudy (anonymous):

Ok Go On geogebra and you will get what Im saying. What Im in in algebra 2 right now is module 7 and the discontinous is (-4, -11). and I just checked it and it was right becasue if you plug in that function on geogebra and you go to the point (-4,-11) it says undefined. You find the value that cant equal zero in this case by substracting 4 from both sides giving you -4. Then you take negative 4 and plug it into the simplified form which was what you told me 2x- 3. therefore, f(-4) = 2(-4) - 3 which gives you the point (-4, -11) which in the discontinuitie of the function and it prooves it when you put it into geogebra becasue it says undefined. get my drift? I just didnt know how to simplfy the numerator on this equation. I need help on simplifying I know how to do everything else.

OpenStudy (anonymous):

ooh i see in that case, since i pretty much suck at factoring, i simply cheat

OpenStudy (anonymous):

i let the computer factor for me http://www.wolframalpha.com/input/?i=2x^2+%2B+5x+-+12 not sure what geoalgbebra is on line class?

OpenStudy (anonymous):

I just didnt and I really still dont understand how to factor something like 2x^2 + 5x - 12 since I cant take 2 from 5

OpenStudy (anonymous):

yeah i know i just write \((2x+...)(x+...)\) and grind it til i find it

OpenStudy (anonymous):

but look, you know from the problem that you are going to end up cancelling with the denominator, so one factor has to be \(x+4\)

OpenStudy (anonymous):

so it is not has hard as it might be put \(2x^2+5x-12)=(x+4)(2x+?)\) and then you have really no choices

OpenStudy (anonymous):

Well I am taking Algebra 2 honors this summer so i dont have to take it next year (10th grade year) so i take it on FLVS, and every teacher recommends geogebra.org . check it out you can do it without even installing or download. when you get on the page click download(it wont) then go to webstart not the geoprim or applet thing go to webstart and it will put it in your browser, but you do have to have java. Dude when I tell you its a lifesaver and teachers RECOMMEND IT, you can just plug in a function with out have to put it in (y=mx+b) you can put it however and it graphs it for you , you can go to object properties and it will tell you the slop and I plug in functions like this for expamply plug in f(x) = (2x^2 + 5x - 12)/(x + 4) press enter and it tells you therzeros and I wish I new about it with geometry but Im glad i have it for Algebra 2. Sorry i just wrote a ton :o ;p;

OpenStudy (anonymous):

i have heard it is good i usually use wolfram because there is nothing to download, and you just type stuff in and get pretty much everything you want

OpenStudy (anonymous):

I just checked that website out and your a lifesaver. Haha Now I wont have to worry about factoring crap like that again :p

OpenStudy (anonymous):

Wolfram makes Mathematica, there is also Maple, by someone else, but GeoGebra looks pretty cool